OpenStudy (anonymous):

A whole number increased by its square is 4 times itself. Find the number.

OpenStudy (anonymous):

so we can approach this question as follows: let's say a whole number is x. so we can say that increased by its squared so x^2 is 4 times itself. so x^2 = 4x therefore... x^2 = 4x and now if we divide by x, then x = 4.

OpenStudy (anonymous):

I disagree that x^2 is a number increased by its square. I'd say that was the number squared.

OpenStudy (anonymous):

I think we're going to end up with a quadratic.

OpenStudy (anonymous):

oh ok. i interpretde it differently. if you talk about increased by its square then you get x + x^2 = 4x in that case you take the x over to the other side ... x^2 = 3x so x = 3